OpenStudy (anonymous):

a ball is thrown upward from a 100-foot tall building with an initial velocity of 14 feet per second, find the interval of time which the ball is greater than 103 feet. s(t)=-16t^+14t+100????

OpenStudy (lgbasallote):

is this the "full" problem? the question did not explain what the equation meant?

OpenStudy (lgbasallote):

it can mean something else sometimes

OpenStudy (anonymous):

it says the height s(t) is given by this function.

OpenStudy (lgbasallote):

ahh better

OpenStudy (lgbasallote):

so when height is 103 feet then s(t) would be 103 do you agree?

OpenStudy (anonymous):

i would think so

OpenStudy (lgbasallote):

well s(t) would be 103 because it says the height is 103 and s(t) is the height. do you get the logic? you need to understand that part because it is crucial ;)

OpenStudy (lgbasallote):

so s(t) = 103 that means \[-16t^2 + 14t + 100 = 103\] \[-16t^2 + 14t + 100 - 103 = 0\] \[-16t^2 + 14t - 3 = 0 \] do you get it so far?

OpenStudy (anonymous):

ok yes i see that

OpenStudy (lgbasallote):

good. now you use the quadratic formula to solve for t. i assume you're familiar with the formula yes?

OpenStudy (anonymous):

somewhat

OpenStudy (lgbasallote):

great. \[t = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] tell me what you get after you solve it :D

OpenStudy (anonymous):

-14 square root 4/32

OpenStudy (lgbasallote):

and what is the square root of 4?

OpenStudy (anonymous):

2

OpenStudy (lgbasallote):

btw...32 should be negative.. 2 x -16 = -32

OpenStudy (lgbasallote):

\[\frac{-14 \pm \sqrt 4}{-32} \implies \frac{-14 \pm 2}{-32}\]

OpenStudy (anonymous):

Thank you so much.

OpenStudy (lgbasallote):

ahh you got the answer already?

OpenStudy (anonymous):

No so i assume it needs to be simplified?

OpenStudy (lgbasallote):

yup.. i can break it down into two parts \[\frac{-14 + 2}{-32}\] and \[\frac{-14 - 2}{-32}\]

OpenStudy (lgbasallote):

solve for those two and you'll get the answer

OpenStudy (anonymous):

ok -3/8 and -1/2 ????

OpenStudy (lgbasallote):

nope..remember negative divided by negative is positive also you cant have negative time

OpenStudy (anonymous):

ahhh 3/8 and 1/2

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

you are awesome thank you so much....

OpenStudy (lgbasallote):

<tips imaginary hat>